OpenStudy (starryknight):

The hypotenuse of a right triangle is 41 in long. One leg of the same triangle is 40 in long. What is the length of the other leg? Explain how you found your answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

So are you familiar with the Pythagorean a^2 = b^2 + c^2? You're given a, b here so you just need to find c.

OpenStudy (anonymous):

See above for the explanation. You can use Pythagorean because it's a right triangle. You're given a, b so you can find c. What else is missing?

OpenStudy (starryknight):

how can you explain this in words (9cm)²+(x)²=(41cm)² 81cm+x²= 1681cm x²= 1600cm x = √1600cm x = 40cm Solution: x = 40cm

OpenStudy (anonymous):

Look at the question again. 40 is given right?

OpenStudy (starryknight):

yeah

OpenStudy (starryknight):

sorry if im annoying you im just bad at math

OpenStudy (anonymous):

Not at all. 40 is given in the question so you don't need to solve for it.

OpenStudy (starryknight):

so i got the problem wrong?

OpenStudy (anonymous):

The hypotenuse of a right triangle is 41 in long. One leg of the same triangle is ***40 in*** long. What is the length of the other leg? Explain how you found your answer. It looks like you're solving above for 40. So is that a typo?

OpenStudy (mertsj):